OpenStudy (anonymous):

lim(x->2 pi) (ceil(cosx))

OpenStudy (anonymous):

Well, \(\cos(2\pi) = 0\), so if you take your right and left hand limits, you're going to have: \[ \Large \lim_{x\rightarrow 2\pi^-}\cos(x) = -1 \]\[\Large \lim_{x\rightarrow 2\pi^+}\cos(x) = 1 \]What does that tell you about the limit?